OpenStudy (evenb0):

I need help solving a system of equations Choose the correct graph of the given system of equations. y − 2x = −1 x + 3y = 4 I'm not looking for the answer, I would like someone to push me in the right direction. Thanks

OpenStudy (evenb0):

I didn't post the options

OpenStudy (khantahmina):

first put each of the equations in y=mx+b form

OpenStudy (evenb0):

would I add y to both sides to accomplish that?

OpenStudy (khantahmina):



OpenStudy (evenb0):

wait

OpenStudy (khantahmina):

y would u wanna do that? just leave the y alone and move the -2x to the other side n u can do this by adding 2x on each side :)

OpenStudy (evenb0):

Now I see

OpenStudy (evenb0):

3y= -x + 4 divide both sides by 3 y= -x+ 1.3 is that right?

OpenStudy (eli_moses):

Remember that when you divide 3 to both sides, the whole equation is divided by 3, including x.

OpenStudy (thatonegirl_):

for the second equation: y=-1/3x+4/3\[y=-\frac{ 1 }{ 3 } x + \frac{ 4 }{ 3 }\]

OpenStudy (thatonegirl_):

Do you understand that part?

OpenStudy (evenb0):

Ehhhhhhh.... No

OpenStudy (evenb0):

I swear this is only because it's math

OpenStudy (thatonegirl_):

No problem, I'll explain. So your starting equation is: \[x + 3y = 4\]

OpenStudy (thatonegirl_):

And you want it to look more normal, meaning you just want a y= bla bla bla

OpenStudy (evenb0):

right

OpenStudy (thatonegirl_):

So to go about that, you will want y by itself. Right?

OpenStudy (evenb0):

subtract x from both sides.

OpenStudy (thatonegirl_):

Good!